what is the no of ways to give at least one rupee to a beggar from a collection of 4 25p. coins,3 50p. coins and 2 Re1 coins?


please answer in detail. The answer is 53 but how to get it?

Grade:10

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
The person has 3 coins of 25 paise, 4 coins of 50 paise and 2 coins of 1 rupee.
The number of ways in which he can give none or some coins to a beggar is (3 + 1)(4 + 1)(2 + 1) = 60 ways.
0 paisa =1
25 paisa=1
50 paisa = 2 ways (1- 50 paisa or 2- 25 paisa)
75 paisa = 2 ways
total = 60-6 = 54
So number of ways to give at least one rupee to a beggar=54